A 0.20 M solution of hydrofluoric acid (HF(aq)) has a pH of 3.96 at 25 oC. Calculate the Ka and Percent Ionization of hydrofluoric acid.

HF(aq) + H2O(l)F-(aq) + H3O+(aq)
1
Expert's answer
2019-08-23T04:15:20-0400

Conditions of question: reaction HF(aq) + H2O(l) = F-(aq) + H3O+(aq)

p-H (solution of hydrofluoric acid) = 3.96 at 25oC

С(HF(aq))=0.20 M

To find: the Ka and Percent Ionization ( α )of hydrofluoric acid  

Solution:  The expression of the acidity constant (Ka) for hydrofluoric acid has the form

"Ka=([F-]*[H30+])\/[HF]"

"pH=-log[H3O+]" , "pKa=-logKa"

Transforming, we get the Henderson-Hasselbach equation:

"-pKa=-pH+log([F-]\/[HF])" or

"pH=pKa+log([F-]\/[HF])".

As stated above, pH=-log[H3O+]. So, [H3O+]=10-3.96=1.097*10-4.

According to the acid dissociation equation [H3O+]=[F-]=1.097*10-4.

Substitute data and get 3,96=pKa + log((1.097*10-4)/0.20)

pKa=3.96-(-3.26)=7.22

Ka=10-7.22=6.062*10-8

Percent Ionization α represents the fraction of disintegrated molecules. If the total concentration of electrolyte (C) is multiplied by this fraction, then we obtain the concentration of any of the ions. So, "Ka=C* \u03b1 *C* \u03b1 \/(C-C* \u03b1 )"

"\u03b1 =" (Ka/C)0.5 = ((6.062*10-8)/0.20)0.5 = 5.489*10-4 = 0.055 %

Answer:  Ka=6.062*10-8, α = 0.055 %.